please help me to solve this equation (x-3)(8x+12): please help me to solve this equation (x-3)(8x+12)
Answer by checkley77(3654)

(x-3)(8x+12)
8x^2+12x-24x-36
8x^2-12x-36
4(2x^2-3x-9)
Question 151504: please help me to solve this equation (x-3)(8x+12): please help me to solve this equation (x-3)(8x+12)
Answer by jim_thompson5910(9403)

This is not an equation. So you cannot solve it, but you can simplify.




(x-3)(8x+12) Start with the given expression.


Now let's FOIL the expression.


Remember, when you FOIL an expression, you follow this procedure:


(highlight(x)-3)(highlight(8x)+12) Multiply the First terms:(x)*(8*x)=8*x^2.


(highlight(x)-3)(8x+highlight(12)) Multiply the Outer terms:(x)*(12)=12*x.


(x+highlight(-3))(highlight(8x)+12) Multiply the Inner terms:(-3)*(8*x)=-24*x.


(x+highlight(-3))(8x+highlight(12)) Multiply the Last terms:(-3)*(12)=-36.


8*x^2+12*x-24*x-36 Now collect every term to make a single expression.


8*x^2-12*x-36 Now combine like terms.


So (x-3)(8x+12) FOILS to 8*x^2-12*x-36.


In other words, (x-3)(8x+12)=8*x^2-12*x-36.